Ticket sales: Tickets for a concert were sold to adults for $3 and to students $2. If the total receipts were $824 and twice as many adult tickets as student tickets were sold, then how many of each were sold ?
Solving this as a linear equation

Answer by checkley75(3666)

The only suggestion I can give you is: Write down every relevent fact and what answer they are looking for. Now state the problem in words then replace the words with whatever numeric values you have and x,y,z,etc for any of the unknown items.
Example: cost of adult tickets 3.00 times the number of adults (2C) is 3*2C.
Now just add the cost of the children's tickets 2.00*C or 2C.
Now the sum of these two costs=824.00 or 3*2C+2C=824.
----------------------------------------------------------------------------
3(2C)+2C=824 $3 FOR EACH ADULT (2*CHILDREN'S). $2 FOR EACH CHILDREN'S TICKET.
6C+2C=824
8C=824
C=824/8
C=103 CHILDREN TICKETS SOLD FOR $2
2*103=206 ADULT TICKETS SOLD.
PROOF
3*206+2*103=824
618+206=824
824=824